In a right triangle c=6 feet 3 inches and tan B=1.2. How do you solve the triangle?
Answer by swincher4391(1107)

It may be best to convert 6ft 3in into 6.25 inches.
Since tan(B) = b/a = 1.2, then b = 1.2a
a^2 + b^2 = (6.25)^2
a^2 + (1.2)^2a^2 = (6.25)^2
a^2 + 1.44a^2 = 39.0625
2.44a^2 = 39.0625
a^2 = 16.0092213
a = 4.000125.
We'll say a = 4.
4^2 + b^2 = (6.25^2)
b^2 = 39.0625 - 16
b^2 = 23.0625
b = 4.80 = 4 ft 9.6 in
Or roughly 4 ft 10 in.
So we have a+=+4, b+=+4.8, c+=+6.25
or a =4, b = 4 ft 10in, c = 6 ft 3 in
I assume we want to find the angles as well?

so A = arcsin(4/6.25) = 39.81º = 40%BA
similarly with B B = arcsin(4.8/6.25) = 50.19º = 50%BA
obviously C is 90º since it is a right triangle.